3. Consider the following information about Nicaragua Year | CPI inflation | Wage | % Real Real change wage wage in wage base 2014 change adjusted real wage ear cordobas wage 2.50 n/a $737.5 | 766.35 199943.98 n/a 2004 | 59.32 | 34.88% | 3.50 | 4090 200995.48 2014135.07 5.687.67 5.907.97 n/a 3.9% 5.50 8.00 The wage is measured as thousands of Nicaraguan Cordobas per month.

a. b. c. d. e. f. Find the inflation rates from 1999-2004, 2004-2009, 2009-2014. Find the percentage change in the wage across the same time periods. Calculate the real wage in each year in base year cordobas. Calculate the real wage in each year in 2014 cordobas. Find the percentage change in the real wage for each time period. Using the real wage values from part d, convert to dollars if the exchange rate is 26 Nicaraguan cordobas per dollar. If goods/services are 60% cheaper in Nicaragua, calculate the purchasing power parity adjusted monthly real wage in each year, using your answers from part f. g.

Need help, please show work. Some correct answers are already shown. All I need to know is how to work through the problems. Please show work
Answer #1

a) Inflation rates from 1999-2004, 2004-2009 & 2009-2014

Inflation rates based on CPI

= CPI of 2004-CPI of 1999/CPI of 1999 x 100

= 59.32-43.98/43.98 x 100

=15.34/43.98 x 100

= 34.879% or 34.88%

= CPI of 2009-CPI of 2004/CPI of 2004 x 100

= 95.48-59.32/59.32 X 100

= 36.16/59.32 x 100

= 60.957% or 60.95%

= CPI of 2014-CPI of 2009/CPI of 2009 x 100

= 135-07-95.48/95.48 X 100

= 39.59/95.48 x 100

= 41.464% or 41.46%

b) Percentage change in wage across years -1999-2004, 2004-2009 & 2009-2014

Percentage change in wage across year: 1999-2004

Wages in 2004-wages in 1999/wages in 1999 x 100

=3.50-2.50/2.50 x 100

= 1/2.50 x 100

= 40%

Wages in 2009-wages in 2004/wages in 2004 x 100

=5.50- 3.50/3.50 x 100

= 2/3.50 x 100

= 57.142 % or 57.14%

Wages in 2014-wages in 2009/wages in 2009 x 100

=8.00- 5.50/5.50 x 100

= 2.5/5.50 x 100

= 45..454 % or 45.45%

c)

Real wage formula: Nominal wage/Price Level x 100

Real wages for year 1999

= 2.50/43.98 x 100

= 5.684 or 5.68

Real wages for year 2004

= 3.50/59.32 x 100

= 5.90

Real wages for year 2009

= 5.50/95.48 x 100

= 5.76

Real wages for year 2014

= 8.00/135.07 x 100

= 5.92

d) Real wage in each year in 2014 cordobas

Real wage in 1999

Real wage = Wages of 1999 x CPI of 2014 / CPI of 1999

= 2.50 x 135.07/43.98

Real wage = 7.67

Real wage in 2004

Real wage = Wages of 2004 x CPI of 2014 / CPI of 2004

= 3.50 x 135.07/59.32

Real wage = 7.969 or 7.67

Real wage in 2009

Real wage = Wages of 2009 x CPI of 2014 / CPI of 2009

= 5.50 x 135.07/ 95.48

Real wage = 7.78

Real wage in 2014

Real wage = Wages of 2014 x CPI of 2014 / CPI of 2014

= 8.00 x 135.07/ 135.07

Real wage = 8